# Orbital magnetization in Sierpinski fractals

### v2

@@ Abstract
Orbital magnetization (OM) in Sierpinski carpet (SC) and triangle (ST) fractal is theoretically investigated by using Haldane model as a prototypical example. The OM calculation is performed following two distinct approaches; employing the definition and local markers formalism. Both methods coincides for all systems analyzed. For the SC, higher fractal generations create a dense set of edge states, resulting in a staircase profile, leading to oscillations in the magnetization as a function of the chemical potential. In contrast, the ST self-similarity produces distinct fractal-induced spectral gaps, which manifest as constant plateaus in the magnetization. The STs exhibit a pronounced sensitivity to edge terminations. Our results reveal how quantum confinement in fractal structures affects the electronic orbital angular momentum, pointing to possible pathways for exploring novel orbitronics in systems with complex geometries.
@@
Introduction: Fractals are mathematical structures exhibiting self-similarity and non-integer dimensions [1]. Fractal structures are often found in nature and appear in contexts such as molecular chemistry [2, 3], biology [4, 5], and geography [6]. Of the many types of fractals, the Sierpinski fractals are some of the most studied. They are constructed from simple geometric shapes by iteratively removing specific regions, resulting in structures that exhibit self-similarity. For instance, the Sierpinski carpet (SC) is constructed through an iterative process of subdividing a solid square into a $3\times 3$ grid of smaller congruent squares and removing the central square. This process is repeated recursively for each of the remaining eight squares, resulting in a structure composed of self-replicated solid squares. An analogous algorithm can be used to generate the Sierpinski triangle (ST) fractal.
@@
In physics, electronic fractals have gained increasing attention in the last few years. With advances in the manipulation of matter at the nanoscale, the fabrication of nanostructured fractal patterns has become possible. Electronic fractal systems at the nanoscale have been constructed using various approaches, including bismuth nanostructures [7], surface-positioned molecules [8], and self-assembled molecular systems [9, 10], where electrons are constrained to move within fractal lattice geometries. The non-integer fractal dimension of this structure enables the emergence of many interesting and exotic phenomena. A remarkable example is the recent proposal in which electronic fractals can host topological phases even without the usual driving terms, such as spin-orbit coupling or magnetic fields, in the electronic Hamiltonian [11]. We also mention the possible loss of correlation between transport coefficients and topological invariants in certain electronic fractals, as pointed out by some numerical studies [12, 13]. Electronic fractals have also been known to exhibit particular charge localization [14, 15] and self-similarity in electronic responses [16, 17, 18, 19]. Furthermore, fractal systems lack translational symmetry, making them interesting platforms for studying local topological aspects of matter [20, 21, 22]. In fact, real-space topology has been studied in fractal systems of both $\mathbb{Z}_{2}$ insulators and Chern insulators, including Sierpinski triangles [23, 24, 25] and Sierpinski carpets [26, 27]. Other types of topological phases in distinct fractal systems have also been explored [28, 29, 30, 31].
@@
Figure 1: (a) Topological phase diagram of the 2D Haldane model. The red mark indicates the parameters used in this work. (b) Four generations of the Sierpinski carpet (SC), with a magnified view of the underlying honeycomb lattice. (c) Energy spectra and density of states for SC G(0-4). The inset shows in-gap states from the topological phase. Yellow shading indicates the Haldane bulk gap for parameters $V_{ab}/t_{2}=3.0$, $\phi=0.7\pi$ [red mark in (a)].
Understanding how magnetism manifests itself in fractal geometries remains an intriguing open question [42], especially in light of advances in nanoscale manipulation and the growing range of materials used in nanoscience. In particular, studying the orbital contribution to the magnetization of fractal nanostructures is appealing, given the strong dependence of the electronic spectrum on the specific fractal geometry. To explore theses effects, we study the Haldane model, a paradigmatic example exhibiting orbital magnetization, on two different Sierpinski fractal architectures. Firstly, we present results for the Sierpinski carpet, focusing on the evolution of equilibrium orbital magnetization of the system, with fractal generation of the Sierpinski carpets. A subsequent analysis for Sierpinski triangles highlights the emergence of fractal-induced energy gaps and the crucial role of edge termination geometry.
Model and Methods: In our study, we consider electrons described by the Haldane Hamiltonianon, which can be written in real space as,
@@
where $t_{1}$ is the intensity of hopping between nearest-neighbor sites $\langle{\bf i},{\bf j}\rangle$ in the underlying honeycomb structure. $V_{\rm ab}$ is the intensity of sublattice potential with, $\tau_{\bf i}=+1(-1)$ for the sites at sublattice $A(B)$. The strength coupling $t_{2}$ between the next-nearest-neighbor sites ($\langle\langle{\bf i},{\bf j}\rangle\rangle$) captures the effect of the staggered flux $\phi$ introduced in the model [43]. In this term, $\nu_{{\bf ij}}=+1(-1)$ for clock (counter clock)-wise orientations of hoppings. The illustrative topological phase diagram of this model in the 2D bulk system is shown in Figure 1 (a). It is composed of two topologically non-trivial phases with Chern numbers $C=\pm 1$ and a topologically trivial phase with zero Chern number ($C=0$). All calculations were performed adopting $t_{2}=t_{1}/3$, $V_{ab}/t_{2}=3.0$, and $\phi=0.7\pi$.
Figure 2: (a) Orbital magnetization vs. chemical potential in the Haldane model for Sierpinski carpet generations G(0-3) (a.1-4) and also for (b) Sierpinski triangle generations G(1-4) (b.1-4). Solid lines represent the results obtained using the definition of orbital magnetization [Eq. (2)], while circles represent the results obtained from the local markers formulation. The shaded yellow region corresponds to 2D bulk gap in the Haldane model with $V_{ab}/t_{2}=3.0$, $\phi=0.7\pi$. Blue areas in panels (b.1-4) mark fractal-induced gaps in $Z$ ST. The insets in panels (a.1-4) show magnified views of orbital magnetization in the 2D bulk gap region for SC.
@@
It is worth noting that, despite this interpretation of the markers, only the macroscopic average of $m({\bf r})$ has physical significance, as discussed in Ref. [47]. In Ref. [48], it was shown that the orbital magnetization obtained from local markers coincides with that obtained from the definition of Eq.(2), provided the macroscopic average is taken over the entire sample area, including both bulk and edge regions.
@@
Figure 3: (a) Illustration of five generations of the Sierpinski Triangle with zigzag edges [$Z$ ST G(0-4)] considered in our calculations. (b) Energy spectra with the corresponding density of states for ZZ ST G(0-4). The inset magnifies the region of in-gap states, which appear in the system subjected to open boundary conditions when parameters are adjusted to reach the topological phase of the model. The yellow shaded region represents the band gap of the 2D bulk Haldane model, obtained with the same coupling constants ($V_{ab}/t_{2}=3.0$ and $\phi=0.7\pi$) used in our calculations within the fractal geometry (red mark in Fig. 1 (a)). The blue shaded strips correspond to the fractal induced energy gaps opened in the energy spectra of ZZ ST.
@@
Sierpinski Carpet: The Sierpinski carpet (SC) is geometrically constructed by systematically inserting vacancies inside a bidimensional region, creating different generations $G(n)$. By increasing the number of vacancies the SC dimension assumes a fractional value of $D=log(8)/log(3)\approx 1.89$ independently of the generation. Adopting a honeycomb lattice mesh inside the carpet, the external lengths are fixed between the generations G(0-3) and the corresponding number of lattice sites are $N_{i}=11772,\ 10512,\ 9456,\ 8688$ sites (i=0,,3)(i=0,...,3), respectively, as shown in gray, blue, red and green in Fig. 1 (b). Both external and internal edges of the carpet are formed by sequences of armchair and zigzag contours. The resulting energy spectrum of SCs G(0-3) are exhibited in panel (c) of Fig. 1 following the previously described colored pattern. The adopting parametrization were adjusted to match with a topological phase ($C=\pm 1$) inside the 2D bulk Haldane diagram with $V_{ab}/t_{2}=3.0$ and $\phi=0.7\pi$, which describes the anomalous quantum Hall effect [see Fig. 1 (a)]. As expected from the bulk results [43], in this phase the finite system subjected to OBC has protected edge states inside the bulk energy gap region. Such states develop a staircase profile as the generation of the fractal is increased [30, 49], which implies in emergent peaks in the total density of states (DOS) for this region, as shown in Fig. 1 (c). In fractals this is characterized by emergent edge states occurring in external and internal edges of the system, the later being related with the fractal generation. For the proposed systems, this pattern results from a competition between the lattice parameters and the fractal order, emerging as the vacancies are inserted, as seen for different types of models and fractals [30].
@@
In Fig. 2 (a.1-4) we present the calculated orbital magnetization, $M$, as a function of the chemical potential, $\mu$, for the first four generations of the Sierpinski carpets, i.e. G(0-3). In each panel, the results of the two distinct computational methods are shown to be in perfect matching; the solid lines represent the magnetization calculated from its fundamental definition (Eq. 2), while the circles are obtained from the local marker with projectors formulation (Eqs. 3 and 8). The shaded yellow region highlights the energy range of the bulk topological gap of the 2D Haldane model. For the initial G(0) generation in panel (a.1), the magnetization exhibits a linear trend within this gap, which is a signature of a finite bulk Chern number. It should be noted that the result in Fig.  2 (a.1), obtained here from the real-space calculation for a flake with OBC [SC G(0)], coincides with those obtained from the $\bf k$}-space calculation in 2D bulk for the same set of parameters [46]. As the fractal generation increases from G(1) to G(3) significant fluctuations appears in panels (a.242-4) and become more frequent. This behavior is a direct consequence of the fractal nature of the system, reflecting the “staircase” profile of the several internal and external edge states that emerge in the energy spectrum, as highlighted in the magnified view of this energy region in the insets of Fig. 2 (a).
@@
Sierpinski Triangle: Similarly to the Sierpinski carpet, the Sierpinski triangle (ST) is constructed, exhibiting a fractional dimension of $D=\log(3)/\log(2)\approx 1.58$. Given the $C_{3}$ rotational symmetry of the triangular structures, by adopting a honeycomb lattice, we have two distinct edge configurations for the full Sierpinski triangle; zigzag (ZZ ST) and armchair (AC ST). As previously reported [16], the electronic properties of ZZ and AC STs are intrinsically distinct. The high sensitivity of ST quantum states, related to its edge configurations, comes directly from the fundamental physics of the quantum confinement in a honeycomb lattice.
We start exploring the ZZ edge configurations, whose generations G(0-4) with the number of sites $N_{0-4}=5773,\ 5030,\ 3924,\ 3125,\ 2450$ are depicted at the top of Fig. 3, in gray, blue, red, green, and orange, respectively, adopting the same parameters within Haldane model of the SCs.
Unlike SCs, the inherent symmetry of STs, promotes the emergence of additional fractal gaps within the energy spectrum, which the principals are highlighted by blue shaded areas in the lower panel of Fig. 3. Within the yellow shaded region of the same figure, a distinct staircase profile is also observed for some of the edge states residing inside the Haldane bulk gap. The states are separated by such induced gaps in both fundamental/excited and Fermi energies. Because of the difference in crystal symmetry compared to the Sierpinski carpets, the electronic properties of the Sierpinski triangle are modified by changing the edge termination, as seen in the results for the zigzag and armchair configurations in Fig. 2(b) and Fig. 4, respectively. Fig. 4 summarizes the electronic and magnetic properties of a 6th generation Sierpinski triangle with armchair edges (AC ST G(6)), as illustrated by its lattice structure in panel (a). The corresponding energy spectrum is presented in panel (b), where the yellow shaded region indicates the Haldane bulk gap, and the blue shaded regions mark the additional energy gaps that arise from the fractal geometry. Panel (c) demonstrates the direct correspondence between the gaped regions of the spectrum and the orbital magnetization. For both types of ST, the fractal-induced gaps produce constant orbital magnetization for different values of the chemical potential, including those within the energy region corresponding to the Haldane 2D bulk gap.
Figure 4: (a) Sierpinski triangle, G(6), armchair edges [AC ST G(6)]. (b) Energy spectrum and density of states for the Haldane model ($V_{ab}/t_{2}=3.0$, $\phi=0.7\pi$). Yellow shading: bulk band gap; blue: fractal-induced gaps. (c) Orbital magnetization versus chemical potential in circles and lines corresponding with definition (Eq. 2), and local markers formulation, respectively, with blue areas indicating energy gaps.

### v3

@@ Abstract
Orbital magnetization (OM) in Sierpinski carpet (SC) and triangle (ST) fractal is theoretically investigated by using Haldane model as a prototypical example. The OM calculation is performed following two distinct approaches; employing the definition and local markers formalism. Both methods coincides for all systems analyzed. For the SC, higher fractal generations create a dense set of edge states, resulting in a staircase profile, leading to fluctuations in the magnetization as a function of the chemical potential. In contrast, the ST self-similarity produces distinct fractal-induced spectral gaps, which manifest as constant plateaus in the magnetization. The STs exhibit a pronounced sensitivity to edge terminations. Our results reveal how quantum confinement in fractal structures affects the electronic orbital angular momentum, pointing to possible pathways for exploring novel orbitronics in systems with complex geometries.
I Introduction
@@
Fractals are mathematical structures exhibiting self-similarity and non-integer dimensions [37]. Fractal structures are often found in nature and appear in contexts such as molecular chemistry [48, 50], biology [52, 59], and geography [24]. Of the many types of fractals, the Sierpinski fractals are some of the most studied. They are constructed from simple geometric shapes by iteratively removing specific regions, resulting in structures that exhibit self-similarity. For instance, the Sierpinski carpet (SC) is constructed through an iterative process of subdividing a solid square into a $3\times 3$ grid of smaller congruent squares and removing the central square. This process is repeated recursively for each of the remaining eight squares, resulting in a structure composed of self-replicated solid squares. An analogous algorithm can be used to generate the Sierpinski triangle (ST) fractal.
@@
In physics, electronic fractals have gained increasing attention in recent years. With advances in the manipulation of matter at the nanoscale, the fabrication of nanostructured fractal patterns has become possible. Electronic nanoscale fractal systems have been constructed using various approaches, including bismuth nanostructures [8], surface-positioned molecules [28], and self-assembled molecular systems [41, 51], where electrons are constrained to move within fractal lattice geometries. The non-integer fractal dimension of this structure enables the emergence of many interesting and exotic phenomena. A remarkable example is the recent proposal in which electronic fractals can host topological phases even without the usual driving terms, such as spin-orbit coupling or magnetic fields, in the electronic Hamiltonian [16]. We also mention the possible loss of correlation between transport coefficients and topological invariants in certain electronic fractals, as pointed out by some numerical studies [23, 18]. Electronic fractals have also been known to exhibit particular charge localization [32, 17] and self-similarity in electronic responses [45, 2, 30, 44]. Furthermore, fractal systems lack translational symmetry, making them interesting platforms for studying local topological aspects of matter [5, 53, 39, 4]. In fact, real-space topology has been studied in fractal systems of both $\mathbb{Z}_{2}$ insulators and Chern insulators, including Sierpinski triangles [42, 7, 43] and Sierpinski carpets [47, 31]. Other types of topological phases have also been explored in distinct fractal systems [38, 11, 33, 9].
@@
Figure 1: (a) Topological phase diagram of the 2D Haldane model. The red marks indicate the parameters used in numerical calculations. Parameters (1) and (2) correspond to $V_{ab}/t_{2}=3.0,8.0$ and $\phi=0.7\pi,0.5\pi$, respectively. (b) Four generations of the Sierpinski carpet (SC), with a magnified view of the underlying honeycomb lattice. (c) Energy spectra and density of states for SC G(0-3), considering point (1) in the phase diagram of panel (a). The inset shows in-gap states from the topological phase. Yellow shading indicates the Haldane bulk energy gap for parameters, $t_{2}=t_{1}/3$, $V_{ab}/t_{2}=3.0$ and $\phi=0.7\pi$ .
Understanding how magnetism manifests itself in fractal geometries remains an intriguing open question [12], especially in light of advances in nanoscale manipulation and the growing range of materials used in nanoscience. In particular, studying the orbital contribution to the magnetization of fractal nanostructures is appealing, given the strong dependence of the electronic spectrum on the specific fractal geometry. To explore these effects, we study the Haldane model, a paradigmatic example exhibiting orbital magnetization, on two different Sierpinski fractal architectures. Firstly, we present results for the Sierpinski carpet, focusing on the evolution of equilibrium orbital magnetization of the system, with fractal generation of the Sierpinski carpets. A subsequent analysis for Sierpinski triangles highlights the emergence of fractal-induced energy gaps and the crucial role of edge termination geometry.
II Model and Methods
In our study, we consider electrons described by the Haldane Hamiltonian, which can be written in real space as,
@@ II Model and Methods
where $t_{1}$ is the intensity of hopping between nearest-neighbor sites $\langle{\bf i},{\bf j}\rangle$ in the underlying honeycomb structure. $V_{\rm ab}$ is the intensity of sublattice potential with, $\tau_{\bf i}=+1(-1)$ for the sites at sublattice $A(B)$. The strength coupling $t_{2}$ between the next-nearest-neighbor sites ($\langle\langle{\bf i},{\bf j}\rangle\rangle$) captures the effect of the staggered flux $\phi$ introduced in the model [21]. In this term, $\nu_{{\bf ij}}=+1(-1)$ for clock (counter clock)-wise orientations of hoppings. The illustrative topological phase diagram of this model in the 2D bulk system is shown in Fig. 1 (a). It is composed of two topologically non-trivial phases with Chern numbers $C=\pm 1$ and a topologically trivial phase with zero Chern number ($C=0$). Unless explicitly mentioned, calculations were performed adopting, $t_{2}=t_{1}/3$, $V_{ab}/t_{2}=3.0$, and $\phi=0.7\pi$.
Figure 2: (a) Orbital magnetization vs. chemical potential in the Haldane model for Sierpinski carpet generations G(0-3) (a.1-4) and also for Sierpinski triangles with (b) zigzag and (c) armchair termination for generations G(1-4) (b.1-4) and (c.1-4), respectively. Solid lines represent the results obtained using the definition of orbital magnetization [Eq. 2], while circles represent the results obtained from the local markers formulation [Eqs.(3-8)]. The shaded yellow region corresponds to 2D bulk gap in the Haldane model with $V_{ab}/t_{2}=3.0$, $\phi=0.7\pi$. Blue areas in panels (b.3-4) and (c.3-4) highlight fractal-induced gaps in ZZ and AC STs. The insets in panels (a.1-4) show magnified views of orbital magnetization in the 2D bulk gap region for SC.
@@
It is worth noting that, despite this interpretation of the markers, only the macroscopic average of $m({\bf r})$ has physical significance, as discussed in Ref. [49]. The nomenclature of each contribution above follows the one established in the derivation based on the Wannier-function approach to the orbital-magnetization formula [55]. The term IC (itinerant current) is associated with the orbital magnetization contribution coming from the net current carried by Wannier functions that circulate through the surface of the solid. The LC term (local circulation) arises from the self-circulation of the Wannier functions. The BC (Berry curvature) term contains the explicit relation between orbital magnetization, Berry curvature, and the Chern number, as one can expect from equations (6) and (7), and makes contact with the Streda formula [6, 54]. In Ref. [58], it was shown that the orbital magnetization obtained from local markers coincides with that obtained from the definition of Eq.(2), provided the macroscopic average is taken over the entire sample area, including both bulk and edge regions.
@@
III Sierpinski Carpet
@@
The Sierpinski carpet (SC) is geometrically constructed by systematically inserting vacancies inside a bidimensional region, creating different generations $G(n)$. By increasing the number of vacancies the SC dimension assumes a fractional value of $D=log(8)/log(3)\approx 1.89$ independently of the generation. Adopting a honeycomb lattice mesh inside the carpet, the external lengths are fixed between the generations G(0-3) and the corresponding number of lattice sites are $N_{i}=11772,\ 10512,\ 9456,\ 8688$ sites (i=0,,3)(i=0,...,3), respectively, as shown in gray, blue, red and green in Fig. 1 (b). Both external and internal edges of the carpet are formed by sequences of armchair and zigzag contours. The resulting energy spectrum of SCs G(0-3) are exhibited in panel (c) of Fig. 1 following the previously described colored pattern. The adopted parametrization was adjusted to match the topological phase ($C=+1$) inside the 2D Haldane’s phase diagram with $V_{ab}/t_{2}=3.0$ and $\phi=0.7\pi$ [point (1) in Fig. 1 (a)], which describes the anomalous quantum Hall effect. As expected from the bulk-edge correspondence [21], in this phase, the finite system subjected to OBC has topologically protected edge states inside the bulk energy gap region [yellow shaded area]. Such states develop a staircase profile as the generation of the fractal is increased [33, 35], which implies in emergent peaks in the total density of states (DOS) for this region, as shown in Fig. 1 (c). In fractals this is characterized by emergent edge states occurring in external and internal edges of the system, the later being related with the fractal generation. For the proposed systems, this pattern results from a competition between the lattice parameters and the fractal pattern, emerging as the vacancies are inserted, as seen for different types of models and fractals [33].
Figure 3: (a) Illustration of several generations of the considered Sierpinski Triangle with armchair edges [AC ST G(1-4)] and (b) with zigzag edges [$Z$ ST G(0-4)]. Energy spectra with the corresponding density of states for AC and ZZ STs. The insets magnify the regions of in-gap states, which appear in the system subjected to open boundary conditions when parameters are adjusted to reach the topological phase of the model. The yellow shaded regions represent the band gaps of the 2D bulk Haldane model, obtained with the same coupling constants ($V_{ab}/t_{2}=3.0$ and $\phi=0.7\pi$) used in our calculations within the fractal geometry [point (1) in Fig. 1 (a)]. The blue shaded strips correspond to the fractal induced energy gaps opened in the energy spectra of ZZ and AC STs.
@@
In Fig. 2 (a.1-4) we present the calculated orbital magnetization, $M$, as a function of the chemical potential, $\mu$, for the first four generations of the Sierpinski carpets, i.e. G(0-3). In each panel, the results of the two distinct computational methods are shown to be in perfect matching; the solid lines represent the magnetization calculated from its fundamental definition (Eq. 2), while the circles are obtained from the local marker with projectors formulation (Eqs. 3 and 8). The shaded yellow region highlights the energy range of the bulk topological gap of the 2D Haldane model. For the initial G(0) generation in panel (a.1), the magnetization exhibits a linear trend within this gap, which is a signature of a finite bulk Chern number. It should be noted that the result in Fig. 2 (a.1), obtained here from the real-space calculation for a flake with OBC [SC G(0)], coincides with those obtained from the $\bf k$}-space calculation in 2D bulk for the same set of parameters [10]. As the fractal generation increases from G(1) to G(3) significant fluctuations appears in panels (a.242-4) and become more frequent. This behavior is a direct consequence of the fractal nature of the system, reflecting the “staircase” profile of the several internal and external edge states that emerge in the energy spectrum, as highlighted in the magnified view of this energy region in the insets of Fig. 2 (a).
Figure 4: Orbital magnetization calculated using local markers formulation (projectors formalism) for (a) ST AC G(4), (b) ST AC G(0) and (c) SC G(3) and its respective contributions for each component $M_{\text{IC}}$ (red dots), $M_{\text{LC}}$ (blue dots), and $M_{\text{BC}}$ (black dots). Haldane bulk gap and fractal induced gaps are highlighted, respectively, by regions in yellow and blue.
IV Sierpinski Triangle
@@
Similarly to the Sierpinski carpet, the Sierpinski triangle (ST) is constructed, exhibiting a fractional dimension of $D=\log(3)/\log(2)\approx 1.58$. Given the $C_{3}$ rotational symmetry of the triangular structures, by adopting a honeycomb lattice, we have two distinct edge configurations for the full Sierpinski triangle; zigzag (ZZ ST) and armchair (AC ST). As previously reported [45], the electronic properties of ZZ and AC STs are intrinsically distinct. The high sensitivity of ST quantum states, related to its edge configurations, comes directly from the fundamental physics of the quantum confinement in a honeycomb lattice.
Fig. 3 shows a comparison of the electronic spectrum of both AC and ZZ edge configurations in panels (a) and (b), respectively. The number of sites in the AC and ZZ ST G(0-4) generations are: $N^{AC}_{0-4}=7056,\ 5400,\ 4212,\ 3402,\ 2916$ and $N^{ZZ}_{0-4}=6721,\ 5279,\ 4307,\ 3737,\ 3506$, illustrated by gray, blue, red, green, and orange, respectively. Since both G(0) AC and ZZ STs have quite similar electronic spectra we have omitted the results regarding the AC ST G(0) energy spectra and DOS. The parameters of the Haldane model are the same adopted for the SCs [point (1) in the phase diagram of Fig. 1 (a)].
Unlike SCs, fractality promotes additional gaps in the energy spectrum of STs. The presence of fractal gaps in STs and their absence in SCs are consistent with previous numerical works [45, 42, 57]. The principal fractal gaps in ST are highlighted by blue shaded areas along the energy states of Fig. 3. Within the yellow shaded region of the same results, a distinct staircase profile is also observed for some of the edge states residing inside the Haldane bulk gap. These states are separated by such fractal-induced gaps. For both types of ST, the fractal-induced gaps produce plateaus with constant orbital magnetization for different values of the chemical potential, as can be observed in Fig. 2 (b.3-4) and (c.3-4). Notably, some of these fractal gaps and orbital magnetization plateaus occur in an energetic region that overlaps with the Haldane bulk gap. In the next section, we further discuss this phenomenon.
V Discussion